# Online Detection of Anomalies in Temporal Knowledge Graphs with Interpretability

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Anomaly detectionccs: Computing methodologies Temporal reasoning
@@ 3.1. Temporal Knowledge Graph
A temporal knowledge graph is denoted as $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T},\mathcal{F})$. $\mathcal{E}$ and $\mathcal{R}$ are entity set and relation set, respectively. $\mathcal{T}$ is the set of observed timestamps and $\mathcal{F}$ is the set of facts. In real-world scenarios, $\mathcal{E}$, $\mathcal{T}$, and $\mathcal{F}$ will be continuously enriched. Each tuple $(s,r,o,t)\in\mathcal{F}$ connects the subject and object entities $s,o\in\mathcal{E}$ via a relation $r\in\mathcal{R}$ in timestamp $t\in\mathcal{T}$, which means a unit knowledge (i.e., a fact). We represent the connectivity of $\mathcal{G}$ in each timestamp $t$ with a $|\mathcal{E}|\times|\mathcal{E}|\times|\mathcal{R}|$ adjacency tensor $A_{t}$, where 1 represents that the entities are connected by the relation in timestamp $t$. There are two common occurring relationships exist in TKGs. One is chain occurring defined as $\{(s,r_{i},o,t_{i})\rightarrow(s,r_{j},o,t_{j})|t_{j}\geq t_{i}\}$, e.g., $(Obama,WintheSelection,UnitedStates,2008/11/04)\rightarrow(Obama$, $Presidentof,UnitedStates,2009/01/20)$. The other is triadic occurring defined as $\{((s,r_{i},o,t_{i}),(s,r_{j},p,t_{j}))\rightarrow(o,r_{k},p,t_{k})|t_{k}\geq max(t_{i},t_{j})\}$, e.g., $((China,HostVisit,SaudiArabia,2023/03/06),\\ (China,HostVisit,Iran,2023/03/06)\rightarrow(SaudiArabia,SignAgree$ $ment,Iran,2023/03/10)$. The facts on the left of the arrow are called head facts, and those on the right are called tail facts.
@@ 3.2. Anomalies in TKGs
Knowledge updating may make existing facts invalid, but update delays will let these invalid facts not be removed from TKGs. Formally, we define the time errors as $\mathcal{F}_{t}=\{(s_{t},r_{t},o_{t},t_{t})|(s_{t},r_{t},o_{t},t_{t})\in\mathcal{F},(s_{t},r_{t},o_{t})\in\hat{\mathcal{L}},(s_{t},r_{t},o_{t},t_{t})\notin\hat{\mathcal{F}}\}$. For example, $(Obama,Presidentof,UnitedStates,2023/10/21)$.
@@ 3.4. Summarization of A TKG
Given a TKG $\mathcal{G}$, we first construct a function $\mathcal{C}(\cdot)$, which takes each entity as input and outputs its category. Based on this, each knowledge $(s,r,o,t)\in\mathcal{G}$ can be mapped as an atomic rule $(\mathcal{C}(s),r,\mathcal{C}(o))$, which summarizes the interaction pattern of the knowledge (e.g., $(Obama,Win,NobelPeacePrize,2019/$ 10/09)10/09)10 / 09 ) can be mapped as $(PERSON,Win,PRIZE)$).
@@ 3.5. Problem Definition
Figure 2. (a) Conceptual illustration of the proposed AnoT framework. (b) Rule graph construction, which contains: 1. Category function construction, 2. Generating candidate atomic rules, 3. Generating candidate rule edges, 4. Ranking and selecting. (c) Conceptual illustrations of the scoring process and the updater module.
@@ 4.1. Overview
In this paper, we propose AnoT, a novel summarization method for TKG anomaly detection. As depicted in Figure 2(a), AnoT takes an online updating TKG as input, identifies anomalies, and then filters valid knowledge. The process initiates with the detector module, which constructs a rule graph based on the offline preserved part of TKG. Upon the arrival of new knowledge, this module evaluates it against the rule graph to compute an anomaly score. Subsequently, the updater module receives valid knowledge identified by the detector module, and then reforms them as edit operations on the rule graph to handle online semantic and pattern changes. The monitor module estimates the approximate error of the rule graph in representing the TKG. When the approximate error exceeds the threshold, the monitor will inform the detector to refresh the rule graph based on the current TKG. In this way, the reachable nodes during walking will give readable evidence for detection, while the complex patterns can be flexibly described by the walking paths, and the online changes are uniformly handled.
@@ 4.2. Description Length of The Rule Graph
(2) $\begin{split}L(M)=&log(2*|\mathcal{C}_{\mathcal{E}}|^{2}*|\mathcal{R}|)+log\binom{2*|\mathcal{C}_{\mathcal{E}}|^{2}*|\mathcal{R}|}{3}\\ &+\sum_{v\in\mathrm{V}}L(v)+\sum_{e\in\mathrm{E}}L(e),\end{split}$
@@ 4.3. Detector
The detector module has two functions: 1) Construct the optimal rule graph. As shown in Figure 2(b), a category function is first constructed based on existing knowledge (Section 4.3.1), and then it will be used to map knowledge as candidate atomic rules and candidate rule edges (Section 4.3.2). Finally, the most expressive candidates will be iteratively selected to construct the rule graph (Section 4.3.3). More details can be found in Algorithm 1. 2) Generate anomaly scores. As shown in the upper part of Figure 2(c), new knowledge will be first mapped as a set of atomic rules. Atomic rules that exist in the rule graph can give evidence of their conceptual validity, which will derive the static scores (Section 4.3.4). Then, the evidence of time validity is gathered by recursively walking on the rule graph and instantiating the processor nodes, which will derive the temporal scores. More details can be found in Algorithm 2.
@@
[…]
@@ 4.4. Updater
The continuous enrichment of TKG will alter the graph structure, entity semantics, and graph patterns, and introduce new entities, requiring ANoT to adapt online. As shown in the bottom part of Figure 2(c), we propose an updater module to flexibly handle these changes. Based on Algorithm 3, we detail the updater as follows.
@@ 4.7. Generalize to Time-duration-based TKGs
Conceptual errors. Time-duration-based and timestamp-based facts only differ in time annotations. Since the detection of conceptual errors only relies on finding conflicts of interaction preference for entities and not related to time information, the proposed static score can be seamlessly used in time-duration-based TKGs.
@@ 5.2. Experimental Setting
(a) ICEWS 14
(b) ICEWS 05-15
(c) YAGO 11k
(d) GDELT
@@
Figure 7. Performance of ANoT under the different settings of the number of entity categories $k$
@@
5.3. RQ1: Overall Evaluation
5.4. RQ2: Effect of Each Component
5.5. RQ3: Efficiency Analysis
5.6. RQ4: Interpretability Analysis
@@ 5.7. RQ5: Generalization Analysis
Effectiveness of different rule graphs. Figure 8(a) shows how four types of rule graphs contribute to the performance. We can see that all four types of rule graphs can describe a unique part of time-duration knowledge, showing their necessities. As the number of entity categories increases, each type of rule graph can describe more knowledge.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: PACMMODjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Computing methodologies Anomaly detectionccs: Computing methodologies Temporal reasoning
@@ 3.1. Temporal Knowledge Graph
[…]
@@ 3.2. Anomalies in TKGs
[…]
@@ 3.4. Summarization of A TKG
Given a TKG $\mathcal{G}$, we first construct a function $\mathcal{C}(\cdot)$, which takes each entity as input and outputs its category. Based on this, each knowledge $(s,r,o,t)\in\mathcal{G}$ can be mapped as an atomic rule $(\mathcal{C}(s),r,\mathcal{C}(o))$, which summarizes the interaction pattern of the knowledge (e.g., $(Obama,Win,NobelPeacePrize,2019/10/09)$ can be mapped as $(PERSON,Win,PRIZE)$).
@@ 4.1. Overview
In this paper, we propose AnoT, a novel summarization method for TKG anomaly detection. As depicted in Figure 2, AnoT takes an online updating TKG as input, identifies anomalies, and then filters valid knowledge. The process initiates with the detector module, which constructs a rule graph based on the offline preserved part of TKG. Upon the arrival of new knowledge, this module evaluates it against the rule graph to compute an anomaly score. Subsequently, the updater module receives valid knowledge identified by the detector module, and then reforms them as edit operations on the rule graph to handle online semantic and pattern changes. The monitor module estimates the approximate error of the rule graph in representing the TKG. When the approximate error exceeds the threshold, the monitor will inform the detector to refresh the rule graph based on the current TKG. In this way, the reachable nodes during walking will give readable evidence for detection, while the complex patterns can be flexibly described by the walking paths, and the online changes are uniformly handled.
Figure 2. Conceptual illustration of the proposed AnoT framework.
@@ 4.2. Description Length of The Rule Graph
(2) $\begin{split}L(M)=log(2*|\mathcal{C}_{\mathcal{E}}|^{2}*|\mathcal{R}|)+log\binom{2*|\mathcal{C}_{\mathcal{E}}|^{2}*|\mathcal{R}|}{3}+\sum_{v\in\mathrm{V}}L(v)+\sum_{e\in\mathrm{E}}L(e),\end{split}$
@@
Figure 3. Rule graph construction, which contains: 1. Category function construction, 2. Generating candidate atomic rules, 3. Generating candidate rule edges, 4. Ranking and selecting.
@@ 4.3. Detector
The detector module has two functions: 1) Construct the optimal rule graph. As shown in Figure 3, a category function is first constructed based on existing knowledge (Section 4.3.1), and then it will be used to map knowledge as candidate atomic rules and candidate rule edges (Section 4.3.2). Finally, the most expressive candidates will be iteratively selected to construct the rule graph (Section 4.3.3). More details can be found in Algorithm 1. 2) Generate anomaly scores. As shown in the upper part of Figure 4, new knowledge will be first mapped as a set of atomic rules. Atomic rules that exist in the rule graph can give evidence of their conceptual validity, which will derive the static scores (Section 4.3.4). Then, the evidence of time validity is gathered by recursively walking on the rule graph and instantiating the processor nodes, which will derive the temporal scores. More details can be found in Algorithm 2.
@@
Figure 4. Conceptual illustrations of the scoring process and the updater module.
@@
Case demonstration. Here we demonstrate how our strategies detect anomalies in Figure 1. $(KimJongUN,HeldMilitaryExercises,SouthKorea,2020/08/23)$ will be assigned with a high $\mathbb{S}$ due to the interaction preference conflict of category $PRESIDENT$, and thus be detected a conceptual error. $(Trump,Presidentof,UnitedStates,2023/01/20)$ will be assigned with a high $\mathbb{T}$ since it has occurrence order conflict with $(Trump,OutgoingPresident,UnitedStates,2021/01/19)$, and thus be a time error. During temporal scoring, ANoT needs to traverse the rule graph to find instantiateable nodes, and $(PERSON,WintheElection,UnitedStates)$ will be found uninstantiateable. By verifying the low $\mathbb{S}$ and $\mathbb{T}$ of instantiating it using $JoeBiden$, this knowledge will be detected as a missing error.
@@ 4.4. Updater
The continuous enrichment of TKG will alter the graph structure, entity semantics, and graph patterns, and introduce new entities, requiring ANoT to adapt online. As shown in the bottom part of Figure 4, we propose an updater module to flexibly handle these changes. Based on Algorithm 3, we detail the updater as follows.
@@ 4.7. Generalize to Time-duration-based TKGs
Conceptual errors. Time-duration-based and timestamp-based facts only differ in their time annotations. Since the detection of conceptual errors only relies on finding the conflicts of interaction preference for entities and not related to time information, the proposed static score can be seamlessly used in time-duration-based TKGs.
@@ 5.3. RQ1: Overall Evaluation
5.4. RQ2: Effect of Each Component
@@
Figure 9. Performance of ANoT under the different settings of the number of entity categories $k$.
5.5. RQ3: Efficiency Analysis
@@ 5.7. RQ5: Generalization Analysis
Effectiveness of different rule graphs. Figure 10(b) shows how four types of rule graphs contribute to the performance. We can see that all four types of rule graphs can describe a unique part of time-duration knowledge, showing their necessities. As the number of entity categories increases, each type of rule graph can describe more knowledge.
@@ 6. Conclusion
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (No. 62276047) and Shenzhen Science and Technology Program (No. JCYJ20210324121213037).